# Scalable DC Optimization via Adaptive Frank-Wolfe Algorithms

### v1

@@ 5.2 Summary of the experiments
The adaptive error bound criterion proves to be highly efficient when used for early stopping (ES). Effectively it dynamically adjusts the subproblem solution precision based on the current DC gap, avoiding unnecessary computational overhead in early iterations. This can be clearly seen in the instance plots’ number of required LMO calls for the subproblems in the early iterations; see Section C.
@@
The ES solvers sometimes exhibit slightly slower convergence in terms of DCA iterations, which aligns with the theoretical prediction in Theorem 3.2, where we trade a factor of 2 in the convergence rate for the removal of additive error terms.
@@
While each of the two strategies ES or WS are very beneficial in terms of reducing computational cost, the combination of the two strategies is the most efficient by a huge margin, enabling multiple orders of magnitude fewer LMO calls and reducing running times by a similar factor as shown in Figure 1. In particular, the combination of the two strategies enables us to tackle instance sizes that would otherwise be out of reach. In fact, in many cases, the improvement is so substantial that we cannot precisely quantify it because the baseline solvers fail to complete within the time limit.
Results across all considered instances (except for QAP) are summarized in Table 1 and Figure 1. The table uses shifted geometric means as customary; more results and examples of representative individual runs are to be found in Section C. We excluded QAP from this overview as the instances typically do not solve to optimality and we thus we report these separately in detail in Section 5.5. As can be seen, the DCA-BPCG-WS-ES variant (combining warm-starting and early stopping) is the most efficient both in time and LMO calls being often by multiple orders of magnitude faster (geometric means) than the other variants. For a very small number of experiments ($\approx 1\%$) it seems that DCA-BPCG-WS outperform DCA-BPCG-WS-ES as can be seen that the very top of the performance profile chart. However this is mostly an artifact of our experimental design and the hard cap on DCA iterations. By Theorem 3.2 we know that the primal progress may be slower in DCA iterations when using the adaptive error bound criterion while each of the DCA iterations is much faster. That is precisely what we see here: these select few experiments finish via DCA-BPCG-WS very close to the DCA iteration limit and the DCA-BPCG-WS-ES variant would require a few extra iterations to close the gap. This can be also seen when compared to the required compute time: DCA-BPCG-WS-ES requires several orders of magnitude less time (geometric mean) than DCA-BPCG-WS so that a few extra iterations are not a problem.
Figure 1: Performance profile over all tested instances (excluding QAP) as described in Section 5.1. The plot demonstrates the superior efficiency of the DCA-BPCG-WS-ES variant (combining warm-starting and early stopping) across the entire range of problem instances.
Table 1: Performance Statistics over all instances (excluding QAP; see Section 5.5) [Shifted Geometric Means, shift: $1.0$].
@@ 5.3 Difference of Convex Quadratic Functions
where $A,B\in\mathbb{R}^{n\times n}$ are positive definite matrices (constructed as $M^{T}M+0.1I$ for random matrices $M$), $a,b\in\mathbb{R}^{n}$ are random vectors, and $c,d\in\mathbb{R}$ are random scalars. The feasible region is the probability simplex $P=\{x\in\mathbb{R}^{n}:x\geq 0,\sum_{i=1}^{n}x_{i}=1\}$. We ran two separate tests, once with smaller to moderate instance sizes (see Table 2 and Figure 2), where we imposed an inner iteration limit of $10000$ iterations and a limit of $200$ DCA iterations, and once with larger instance sizes (see Table 3 and Figure 3 and Figure 3), where we imposed an inner iteration limit of $50000$ iterations and a limit of $500$ DCA iterations. In both cases, we set the DCA gap tolerance to $10^{-6}$ and the Frank-Wolfe gap tolerance to $210^{-6}/$ 2.
@@
Figure 2: Performance profile over small to medium-sized instances of differences of convex quadratics from Section 5.3.
Table 2: Performance Statistics over small to medium-sized instances of differences of convex quadratics from Section 5.3 [Shifted Geometric Means, shift: $1.0$].
@@
Figure 3: Performance profile over medium to large-sized instances of differences of convex quadratics from Section 5.3.
Table 3: Performance Statistics over medium to large-sized instances of differences of convex quadratics from Section 5.3 [Shifted Geometric Means, shift: $1.0$].
@@ 5.4 Hard DC function example
over the $k$-sparse polytope $C_{\tau,k}=B_{1}(k\tau)\cap B_{\infty}(\tau)=\{x\in\mathbb{R}^{n}:\|x\|_{1}\leq k\tau,\|x\|_{\infty}\leq\tau\}$ with $\tau=10$ and $k=10$, which is the intersection of the $\ell_{1}$-ball of radius $k\tau$ and the $\ell_{\infty}$-ball of radius $\tau$. This example is particularly challenging because the exponential term in $f(x)$ can lead to rapid growth, while the logistic terms in $g(x)$ introduce additional nonlinearity. The $k$-sparse polytope further complicates the optimization landscape. The matrices $A$, $B$ and the vectors $a,b,c,d\in\mathbb{R}^{n}$ are constructed as in Section 5.3.
Results are reported in Table 4 and Figure 4 and match the observations from Section 5.3.
Figure 4: Performance profile over hard DC instances from Section 5.4.
Table 4: Performance Statistics for hard DC instances from Section 5.4 [Shifted Geometric Means, shift: $1.0$].
@@ 5.5 Quadratic Assignment Problem
While projection onto the Birkhoff polytope is computationally expensive, the LMO can be implemented efficiently using the Hungarian algorithm, which is readily available in the FrankWolfe.jl package. We evaluate the algorithms on instances from QAPLIB [Burkard et al., 1997], a well-established benchmark library for quadratic assignment problems111available at https://coral.ise.lehigh.edu/data-sets/qaplib/qaplib-problem-instances-and-solutions/. QAPLIB contains a wide range of QAP instances that have been used extensively in the literature to benchmark QAP algorithms. From the 136 instances in QAPLIB, we successfully processed 97 instances (71.3%) for our experiments. The remaining 39 instances (28.7%) could not be processed due to various data format inconsistencies or parsing issues; for details see Section A.
For our experiments, we used a DCA iteration limit of $500$ and an inner iteration limit for the Frank-Wolfe iterations of $50000$. The instance characteristics can be found in Table 5. Note that the actual problem dimension is quadratic in $n$. Due to their considerable size and the substantially higher LMO complexity (as we have to now run the Hungarian algorithm compared to e.g., sorting before), these instances are much harder than the previous examples, also often coming with objective values that are orders of magnitude larger, so that the additive tolerance is relatively tighter. Thus, the target DCA tolerance of $10^{-6}$ could not always be achieved, even for the fastest algorithm configuration.
Given these computational challenges, we present two complementary performance profiles for these experiments: the classical performance profile in Figure 5, and a modified performance profile in Figure 6 where we consider all instances to be “solved” at their final iteration. This second profile allows us to compare the computational cost—both in terms of runtime and LMO calls—required to reach the DCA iteration limit across different algorithm variants.
Also note that in Table 5, the results seem to be a little bit more noisy, which in fact is due to often having only one QAP instance in a given size bracket. Due to nonconvexity, the optimization paths can differ signficantly, so that sometimes the slower algorithms can be faster as they converged to a different stationary point.
Figure 5: Performance profile over Quadratic Assignment Problem instances from Section 5.5.
Figure 6: Modified performance profile over Quadratic Assignment Problem instances from Section 5.5, where we consider all instances to be solved in their last iteration.
Table 5: Performance Statistics Quadratic Assignment Problem instances from Section 5.5 [Shifted Geometric Means, shift: $1.0$].
@@ B Boosted DCA
In this section we present the preliminary results of the boosted DCA variant. We implemented a simple two level line search via refined grids to compute the step size; relatively exact but not the most efficient. The results are presented in Figure 7. We can see that the boosted DCA variant is basically identical to the unboosted DCA-BPCG-WS-ES variant. Neither is there a significant improvement in DCA iterations nor is it much slower in time due to the additional line search. Note however, that these are preliminary results and an in-depth analysis is left for future work.
@@ C.1 Difference of Convex Quadratic Functions
In this section we present runs of some selected Difference of Convex Quadratic Functions examples from Section 5.3 in Figures 8, 9, 10 and 11; we took the largest sizes for presentation here as the results are more interesting.
@@ C.2 Hard DC function example
In this section we present runs of the hard DC function example from Section 5.4 in Figures 12, 13, 14 and 15.
@@ C.3 Quadratic Assignment Problem
In this section we present runs of some selected QAP instances from Section A in Figures 16, 17, 18 and 19.
Figure 8: Run of the different variants on a Difference of Convex Quadratics example with seed $a4fd7a7efdf1f7d1$ with size $n=1000$ from Section 5.3.
Figure 9: Run of the different variants on a Difference of Convex Quadratics example with seed $e337c88623524354$ with size $n=1000$ from Section 5.3.
Figure 10: Run of the different variants on a Difference of Convex Quadratics example with seed $f308059bfbdf1ef$ with size $n=1000$ from Section 5.3.
Figure 11: Run of the different variants on a Difference of Convex Quadratics example with seed $5931de02be2505f$ 593 with size $n=500$ from Section 5.3.
Figure 12: Run of the different variants on the hard DC function example with size $n=400$ from Section 5.4.
Figure 13: Run of the different variants on the hard DC function example with size $n=500$ from Section 5.4.
Figure 14: Run of the different variants on the hard DC function example with size $n=300$ from Section 5.4.
Figure 15: Run of the different variants on the hard DC function example with size $n=150$ from Section 5.4.
Figure 16: Run of the different variants on the small-sizedesc32b QAP instance from Section A.
Figure 17: Run of the different variants on the medium-sized tai60b QAP instance from Section A.
Figure 18: Run of the different variants on the medium-sized tai64c QAP instance from Section A. For this instance the DCA gaps are oscillating wildly.
Figure 19: Run of the different variants on the large-sized tai100a QAP instance from Section A.

### v2

@@ 5.2 Summary of the experiments
The adaptive error bound criterion proves to be highly efficient when used for early stopping (ES). Effectively it dynamically adjusts the subproblem solution precision based on the current DC gap, avoiding unnecessary computational overhead in early iterations. This can be clearly seen in the instance plots’ number of required LMO calls for the subproblems in the early iterations; see Section˜C.
@@
The ES solvers sometimes exhibit slightly slower convergence in terms of DCA iterations, which aligns with the theoretical prediction in Theorem˜3.2, where we trade a factor of 2 in the convergence rate for the removal of additive error terms.
@@
While each of the two strategies ES or WS are very beneficial in terms of reducing computational cost, the combination of the two strategies is the most efficient by a huge margin, enabling multiple orders of magnitude fewer LMO calls and reducing running times by a similar factor as shown in Figure˜1. In particular, the combination of the two strategies enables us to tackle instance sizes that would otherwise be out of reach. In fact, in many cases, the improvement is so substantial that we cannot precisely quantify it because the baseline solvers fail to complete within the time limit.
[…]
Figure 1: Performance profile over all tested instances (excluding QAP) as described in Section˜5.1. The plot demonstrates the superior efficiency of the DCA-BPCG-WS-ES variant (combining warm-starting and early stopping) across the entire range of problem instances.
Table 1: Performance Statistics over all instances (excluding QAP; see Section˜5.5) [Shifted Geometric Means, shift: 1.01.01.0].
@@ 5.3 Difference of Convex Quadratic Functions
where $A,B\in\mathbb{R}^{n\times n}$ are positive definite matrices (constructed as $M^{T}M+0.1I$ for random matrices $M$), $a,b\in\mathbb{R}^{n}$ are random vectors, and $c,d\in\mathbb{R}$ are random scalars. The feasible region is the probability simplex $P=\{x\in\mathbb{R}^{n}:x\geq 0,\sum_{i=1}^{n}x_{i}=1\}$. We ran two separate tests, once with smaller to moderate instance sizes (see Table˜2 and Figure˜2), where we imposed an inner iteration limit of 100001000010000 iterations and a limit of 200200200 DCA iterations, and once with larger instance sizes (see Table˜3 and Figure˜3 and Figure˜3), where we imposed an inner iteration limit of 500005000050000 iterations and a limit of 500500500 DCA iterations. In both cases, we set the DCA gap tolerance to $10^{-6}$ and the Frank-Wolfe gap tolerance to $10^{-6}/2$.
@@
Figure 2: Performance profile over small to medium-sized instances of differences of convex quadratics from Section˜5.3.
Table 2: Performance Statistics over small to medium-sized instances of differences of convex quadratics from Section˜5.3 [Shifted Geometric Means, shift: 1.01.01.0].
@@
Figure 3: Performance profile over medium to large-sized instances of differences of convex quadratics from Section˜5.3.
Table 3: Performance Statistics over medium to large-sized instances of differences of convex quadratics from Section˜5.3 [Shifted Geometric Means, shift: 1.01.01.0].
@@ 5.4 Hard DC function example
over the $k$-sparse polytope $C_{\tau,k}=B_{1}(k\tau)\cap B_{\infty}(\tau)=\{x\in\mathbb{R}^{n}:\|x\|_{1}\leq k\tau,\|x\|_{\infty}\leq\tau\}$ with $\tau=10$ and $k=10$, which is the intersection of the $\ell_{1}$-ball of radius $k\tau$ and the $\ell_{\infty}$-ball of radius $\tau$. This example is particularly challenging because the exponential term in $f(x)$ can lead to rapid growth, while the logistic terms in $g(x)$ introduce additional nonlinearity. The $k$-sparse polytope further complicates the optimization landscape. The matrices $A$, $B$ and the vectors $a,b,c,d\in\mathbb{R}^{n}$ are constructed as in Section˜5.3.
Results are reported in Table˜4 and Figure˜4 and match the observations from Section˜5.3.
Figure 4: Performance profile over hard DC instances from Section˜5.4.
Table 4: Performance Statistics for hard DC instances from Section˜5.4 [Shifted Geometric Means, shift: 1.01.01.0].
@@ 5.5 Quadratic Assignment Problem
While projection onto the Birkhoff polytope is computationally expensive, the LMO can be implemented efficiently using the Hungarian algorithm, which is readily available in the FrankWolfe.jl package. We evaluate the algorithms on instances from QAPLIB [Burkard et al., 1997], a well-established benchmark library for quadratic assignment problems111available at https://coral.ise.lehigh.edu/data-sets/qaplib/qaplib-problem-instances-and-solutions/. QAPLIB contains a wide range of QAP instances that have been used extensively in the literature to benchmark QAP algorithms. From the 136 instances in QAPLIB, we successfully processed 97 instances (71.3%) for our experiments. The remaining 39 instances (28.7%) could not be processed due to various data format inconsistencies or parsing issues; for details see Section˜A.
For our experiments, we used a DCA iteration limit of 500500500 and an inner iteration limit for the Frank-Wolfe iterations of 500005000050000. The instance characteristics can be found in Table˜5. Note that the actual problem dimension is quadratic in $n$. Due to their considerable size and the substantially higher LMO complexity (as we have to now run the Hungarian algorithm compared to e.g., sorting before), these instances are much harder than the previous examples, also often coming with objective values that are orders of magnitude larger, so that the additive tolerance is relatively tighter. Thus, the target DCA tolerance of $10^{-6}$ could not always be achieved, even for the fastest algorithm configuration.
Given these computational challenges, we present two complementary performance profiles for these experiments: the classical performance profile in Figure˜5, and a modified performance profile in Figure˜6 where we consider all instances to be “solved” at their final iteration. This second profile allows us to compare the computational cost—both in terms of runtime and LMO calls—required to reach the DCA iteration limit across different algorithm variants.
Also note that in Table˜5, the results seem to be a little bit more noisy, which in fact is due to often having only one QAP instance in a given size bracket. Due to nonconvexity, the optimization paths can differ signficantly, so that sometimes the slower algorithms can be faster as they converged to a different stationary point.
Figure 5: Performance profile over Quadratic Assignment Problem instances from Section˜5.5.
Figure 6: Modified performance profile over Quadratic Assignment Problem instances from Section˜5.5, where we consider all instances to be solved in their last iteration.
Table 5: Performance Statistics Quadratic Assignment Problem instances from Section˜5.5 [Shifted Geometric Means, shift: 1.01.01.0].
@@ B Boosted DCA
In this section we present the preliminary results of the boosted DCA variant. We implemented a simple two level line search via refined grids to compute the step size; relatively exact but not the most efficient. The results are presented in Figure˜7. We can see that the boosted DCA variant is basically identical to the unboosted DCA-BPCG-WS-ES variant. Neither is there a significant improvement in DCA iterations nor is it much slower in time due to the additional line search. Note however, that these are preliminary results and an in-depth analysis is left for future work.
@@ C.1 Difference of Convex Quadratic Functions
In this section we present runs of some selected Difference of Convex Quadratic Functions examples from Section˜5.3 in Figures˜8, 9, 10 and 11; we took the largest sizes for presentation here as the results are more interesting.
@@ C.2 Hard DC function example
In this section we present runs of the hard DC function example from Section˜5.4 in Figures˜12, 13, 14 and 15.
@@ C.3 Quadratic Assignment Problem
In this section we present runs of some selected QAP instances from Section˜A in Figures˜16, 17, 18 and 19.
Figure 8: Run of the different variants on a Difference of Convex Quadratics example with seed $a4fd7a7efdf1f7d1$ with size $n=1000$ from Section˜5.3.
Figure 9: Run of the different variants on a Difference of Convex Quadratics example with seed $e337c88623524354$ with size $n=1000$ from Section˜5.3.
Figure 10: Run of the different variants on a Difference of Convex Quadratics example with seed $f308059bfbdf1ef$ with size $n=1000$ from Section˜5.3.
Figure 11: Run of the different variants on a Difference of Convex Quadratics example with seed $1de02be2505f593$ with size $n=500$ from Section˜5.3.
Figure 12: Run of the different variants on the hard DC function example with size $n=400$ from Section˜5.4.
Figure 13: Run of the different variants on the hard DC function example with size $n=500$ from Section˜5.4.
Figure 14: Run of the different variants on the hard DC function example with size $n=300$ from Section˜5.4.
Figure 15: Run of the different variants on the hard DC function example with size $n=150$ from Section˜5.4.
Figure 16: Run of the different variants on the small-sizedesc32b QAP instance from Section˜A.
Figure 17: Run of the different variants on the medium-sized tai60b QAP instance from Section˜A.
Figure 18: Run of the different variants on the medium-sized tai64c QAP instance from Section˜A. For this instance the DCA gaps are oscillating wildly.
Figure 19: Run of the different variants on the large-sized tai100a QAP instance from Section˜A.